OpenStudy (anonymous):

you invest 2,000$ at a rate 4% apr for a 3 months? how much did you earn?

OpenStudy (anonymous):

\[2000*.04*3=?\]

OpenStudy (anonymous):

I = PRT I = 2000 * 0.04 * 3/12

OpenStudy (anonymous):

Helpme...don't u need the time in terms of a year?

OpenStudy (anonymous):

how did you get the 3/12

OpenStudy (anonymous):

Times is always in terms of 1 year. 3 months = 3/12 of a year

OpenStudy (agent0smith):

Because the percentage rate is per year. 3 months is 3/12 of one year.

OpenStudy (anonymous):

What did you get?

OpenStudy (anonymous):

20

OpenStudy (anonymous):

yes $20

OpenStudy (anonymous):

thank you